A Book company is facing following demand function:

Qx = 12,000 – 5,000Px + 5I + 500Pc

where Px = Rs.5,  I = Rs.10,000,   Pc = Rs.6

  1.Determine what effect a price increase would have on  

    total revenue?

 2. Evaluate how sale of books would change during a  

    period of rising incomes.



1
Expert's answer
2021-03-29T11:41:56-0400
"TR=p_xQ_x"

"TR=(12,000-5,000p_x+5I+500p_c)p_x"

"\frac {\delta TR}{\delta p_x}=12,000-10,000p_x+5I+500p_c"

"\frac {\delta TR}{\delta p_x}=12,000-10,000\times 0.5+5\times 10,000+500\times 0.6>0"

TR will increasing when price will increasing.

As income rises, book sales will increase as income enters the function with a positive coefficient.